OpenStudy (anonymous):

Find the indicated root? 3√343

OpenStudy (perl):



OpenStudy (perl):

is there a number you can cube to get 343

OpenStudy (perl):

why dont you start by making a list 1^3 = 1 2^3 = 8 3^3 = 27 4^3 =

OpenStudy (anonymous):

64?

Miracrown (miracrown):

cube root 7

Miracrown (miracrown):